# MGPC: Multimodal Network for Generalizable Point Cloud Completion With Modality Dropout and Progressive Decoding

### v1

@@ I Introduction
As a widely used 3D representation [36], point clouds play a crucial role in vision, computer graphics, augmented reality, and robotics. Despite their popularity, point clouds captured by modern 3D sensors, such as depth cameras and LIDAR scanners, are often incomplete and sparse due to limited viewpoints, occlusions, and measurement noise [13]. Recovering complete object geometry from partial observations is therefore a fundamental problem and is critical for many downstream applications, including 3D reconstruction [47], manipulation [21], and scene understanding [67].
With the rapid advancement of deep learning, point cloud completion has been increasingly studied in a data-driven fashion. Early approaches process voxelized representations using 3D CNNs [18, 6, 17, 63], which suffer from high computational cost and limited scalability as resolution increases. Moving beyond voxel grids, PointNet [36] and its variants [37, 49, 50] inspired subsequent works [71, 48, 2, 24, 72, 51] to directly operate on unordered point sets through symmetric aggregation functions such as max-pooling. More recently, Transformer-based architectures [69, 22, 60, 77] have been introduced to better capture long-range dependencies within point clouds. Although these methods achieve strong performance on existing synthetic benchmarks, their generalization ability to novel objects and real-world scenarios remains limited [43, 54, 32]. As shown in Fig. 1(b), state-of-the-art models trained on existing datasets often produce incomplete shapes and scattered outliers on a previously unseen real-world object.
@@
Through systematic analysis, we identify three key factors that hinder the generalization of existing point cloud completion methods: (1) Limited modality. Most prior works [71, 69, 77] rely solely on partial point clouds as input. However, geometry alone may be insufficient to resolve shape ambiguity, as distinct objects can exhibit similar geometric structures under certain viewpoints. Incorporating complementary modalities, such as RGB images and textual descriptions, provides additional appearance and semantic cues that are beneficial for robust generalization. (2) Limited scalability. Due to the relatively small scale of existing datasets, many methods depend on strong inductive biases, including local operators (e.g., CNNs [18, 6, 17, 63] and KNN-based modules [69, 22, 60, 77]), handcrafted shape priors [66, 5, 11, 59], or carefully engineered network components [3, 40, 73], as summarized in the first row of Fig. 1(a). While effective in low-data regimes, such designs may constrain scalability and reduce the gains achievable as data scale increases [25, 52, 19, 57]. (3) Limited generative capacity. The generative capacity of the generator largely determines the fidelity of completed point clouds. Many existing methods adopt folding-based generators [65, 71, 69] or MLP-based generators [70, 73], which deform a 2D grid into 3D surfaces or regress coordinates in a one-shot manner. Such parameterizations often have limited expressiveness for modeling complex 3D geometric distributions, which can hinder fine-detail recovery and lead to suboptimal reconstruction quality.
@@ II-A Single-Modal Point Cloud Completion
The majority of existing point cloud completion methods take partial point clouds as the sole input modality [71, 69, 77]. These works can be broadly categorized into 3D CNN-based methods [18, 6, 17, 63], point-based methods [71, 48, 2, 24, 72, 51], and Transformer-based methods [69, 22, 60, 77].
@@
Early approaches [18, 6, 17, 63] voxelize point clouds into regular grids and apply 3D CNNs in a manner analogous to 2D image processing. While effective at low resolutions, voxel-based representations incur prohibitive computational and memory costs as resolution increases, which severely limits their scalability. Motivated by the success of PointNet [36] and its variants [37, 49, 50], PCN [71] and subsequent point-based methods [48, 2, 24, 72, 51] operate directly on unordered point sets via shared MLPs, typically coupled with folding-based decoders [65]. For example, CRN [48] introduces a two-stage pipeline that refines coarse predictions through a lifting module, while ODGNet [2] employs an orthogonal dictionary to guide training. However, both approaches require learning dataset-specific shape priors, which may restrict generalization to unseen objects. MSN [24] and NSFA [72] formulate point cloud generation as a deformation process, whereas TopNet [42] and PMPNet [51] model the generation process using tree structures and earth mover formulations, respectively. Despite their effectiveness, many of these methods depend on tightly constrained generators with limited representational capacity, which can lead to oversmoothed surfaces and suboptimal reconstruction quality.
@@
Inspired by advances in natural language processing and 2D vision, recent studies have introduced Transformer architectures [45] to capture global dependencies within point clouds. As a pioneering work, PoinTr [69] proposes a geometric-aware Transformer that incorporates KNN operators to exploit local geometric priors. Subsequent works largely follow two main directions: introducing additional inductive biases [66, 5, 11, 59] and designing specialized architectural components [3, 40, 73]. In the first line, PPCC [66] and PaCo [5] investigate completion based on object parts and planar primitives, respectively. TCorresNet [11] and SPACNet [59] resort to external data structures such as spherical templates and interfaces, which often require extensive pre-processing and hyperparameter tuning. In the second line, SDT [73] and PointAttn [46] study local-global fusion strategies, while SeedFormer [77] and ProxyFormer [22] focus on modifications of Transformer architectures tailored to missing-part sensitivity and upsampling. Despite these efforts, the increasing reliance on complex inductive biases and intricate module designs may limit scalability and generalization under large-scale data regimes [57, 19]. In contrast, we shift our focus to more scalable design based on pure self-attention and cross-attention, combined with large-scale training [25, 52, 19] and further exploit complementary information from multiple modalities to improve completion quality.
@@ II-B Cross-Modal Point Cloud Completion
Cross-modal point cloud completion methods include multi-view approaches [78, 14, 62, 61, 53] and image-assisted approaches [74, 1, 79, 55, 64, 20, 33]. Multi-view methods, such as SVDFormer [78] and AEDNet [14], follow a common paradigm of projecting partial point clouds into depth images from multiple viewpoints, which are then used as auxiliary inputs. SparseNet [62] and MAL-SPC [53] further introduce rendering-based losses by comparing projections of the completed point clouds with Ground-Truth (GT) projections. However, in these approaches, the image modality is derived directly from the input point cloud itself, and thus does not provide additional external information beyond the original geometry.
Another line of research incorporates RGB images to assist point cloud completion [74, 1, 79, 55, 64, 20, 33]. ViPC [74] addresses cross-modal learning by training a generative network to regress complete point clouds from single-view images. Nevertheless, bridging RGB apperance to precise 3D geometry is non-trivial, which may limit ViPC in recovering accurate and complete geometry, especially for occluded regions. Subsequent methods, such as XMFNet [1] and CSDN [79], improve cross-modal fusion in the latent space using cross-attention mechanisms and affine normalization, respectively. EGIINet [64] introduces a feature transfer loss to explicitly guide the modality alignment. More recently, BfR [20] explores a retrieval-based paradigm that leverages similar reference samples from a database. However, its performance heavily depends on the quality and scale of the database and incurs additional computational overhead during inference. In contrast to existing approaches that primarily focus on sophisticated modality fusion strategies, we adopt a unified framework that integrates point clouds, RGB images, and textual descriptions for point cloud completion. Furthermore, the proposed modality dropout mechanism effectively reduces overfitting while improving robustness and flexibility under missing-modality scenarios.
@@ II-C Point Cloud Completion Datasets
Dataset plays a crucial role for learning-based point completion methods [25]. Early efforts [31] leverage 3D models from the ModelNet [58] dataset to evaluate shape completion. PCN [71] subsequently introduces the first widely adopted benchmark based on high-quality ShapeNet [4] models. Later datasets expand upon PCN by incorporating more categories and viewpoints. In particular, ShapeNet-34 and ShapeNet-55 [69] utilize all 55 categories from ShapeNetCore [4] and stratify test samples according to different difficulty levels. However, the crop-based strategy used to generate partial point clouds deviates from realistic sensing processes of depth cameras and laser scanners. The MVP dataset [35] addresses this limitation by rendering partial observations using diverse and uniformly distributed camera poses, although the number of paired samples remains relatively limited. KITTI [15] is frequently employed to evaluate real-world generalization, but it contains only vehicle point clouds and lacks ground-truth complete shapes. More recently, ViPC [74] constructs a large-scale dataset with paired RGB images to facilitate cross-modal learning.
@@ III-D Progressive Decoding
The generator maps fused features back to 3D coordinates, which directly impacts the quality of completed point clouds [30]. Previous methods commonly adopt folding-based or MLP-based generators [71, 69, 70], which either deform a 2D grid into 3D surfaces or regress coordinates in a single step. While effective in simple settings, these generators often have limited generative capacity to represent complex and multimodal geometric distributions, and the one-shot regression paradigm can be sensitive to missing observations and prone to producing uneven point densities. Inspired by point cloud upsampling literature [38, 10, 68], we design a progressive generator that performs coarse-to-fine reconstruction through iterative refinement.
@@ IV-B MGPC-1M Benchmark
As summarized in Table I, the training split of MGPC-1M covers over 1k categories spanning a broad range of daily objects, offering substantially greater category diversity than existing point cloud completion datasets. MGPC-1M contains over one million training pairs, exceeding the combined scale of prior benchmarks in this domain. Unlike most existing datasets that provide only point clouds, MGPC-1M additionally includes RGB images and text descriptions, supporting multimodal evaluation and learning. Beyond scale, MGPC-1M adopts more realistic data generation protocols, including depth back-projection with sensor-like noise and input-centric normalization based on the partial observation, which better matches real-world deployment where ground-truth geometry is unavailable. In contrast to datasets that rely solely on synthetic data, our test split is constructed from the real-world GSO dataset and is entirely disjoint from the training data, enabling rigorous evaluation of cross-domain generalization [28, 19, 27]. The resulting test set contains 1,031 novel instances with high-quality real-world models.
@@ VI Conclusion
In this paper, we present a multimodal framework for generalizable point cloud completion that jointly leverages geometric and semantic cues from point clouds, RGB images, and text within a unified Transformer-based architecture. A modality dropout strategy is introduced to improve robustness under varying modality availability, and a progressive decoder is designed to enable coarse-to-fine reconstruction with improved structural consistency and visual fidelity. To facilitate scalable learning and evaluation, we further develop a VLM-assisted data generation pipeline and construct a large-scale benchmark covering over 1,000 categories and one million training pairs. Extensive experiments on both the benchmark and in-the-wild real-world data demonstrate that the proposed approach consistently improves reconstruction quality and exhibits strong zero-shot generalization, supporting practical 3D perception and visualization applications.

### v2

@@ I Introduction
As a widely used 3D representation [18], point clouds play a crucial role in spatial perception and measurement systems. Despite their popularity, point clouds captured by modern 3D sensors, such as depth cameras and LIDAR scanners, are often incomplete and sparse due to limited viewpoints, occlusions, and inherent sensing uncertainties [15]. Recovering complete object geometry from partial observations is therefore a fundamental problem and is critical for many downstream applications, including pose estimation [44], manipulation [22], and industrial inspection [46].
With the rapid advancement of deep learning, point cloud completion has been increasingly studied in a data-driven fashion. Early approaches process voxelized representations using 3D CNNs [6, 56, 67], which suffer from high computational cost and limited scalability as resolution increases. Moving beyond voxel grids, PointNet [32] and its variants [33, 47, 8] inspired subsequent works [62, 45, 2, 26, 63, 48] to directly operate on unordered point sets through symmetric aggregation functions such as max-pooling. More recently, Transformer-based architectures [60, 23, 53, 66] have been introduced to better capture long-range dependencies within point clouds. Although these methods achieve strong performance on existing synthetic benchmarks, their generalization ability to novel objects and real-world scenarios remains limited [39, 25]. As shown in Fig. 1(b), state-of-the-art models trained on existing datasets often produce incomplete shapes and scattered outliers on a previously unseen real-world object.
@@
Through systematic analysis, we identify three key factors that hinder the generalization of existing point cloud completion methods: (1) Limited modality. Most prior works [62, 60, 66] rely solely on partial point clouds as input. However, geometry alone may be insufficient to resolve shape ambiguity, as distinct objects can exhibit similar geometric structures under certain viewpoints. Incorporating complementary modalities, such as RGB images and textual descriptions, provides additional appearance and semantic cues that are beneficial for robust generalization. (2) Limited scalability. Due to the relatively small scale of existing datasets, many methods depend on strong inductive biases, including local operators (e.g., CNNs [6, 56, 67] and KNN-based modules [60, 23, 53, 66]), handcrafted shape priors [59, 5, 12, 50], or carefully engineered network components [3, 21, 35], as summarized in the first row of Fig. 1(a). While effective in low-data regimes, such designs may constrain scalability and reduce the gains achievable as data scale increases [27, 49, 19, 51]. (3) Limited generative capacity. The generative capacity of the generator largely determines the fidelity of completed point clouds. Many existing methods adopt folding-based generators [58, 62, 60] or MLP-based generators [61, 21], which deform a 2D grid into 3D surfaces or regress coordinates in a one-shot manner. Such parameterizations often have limited expressiveness for modeling complex 3D geometric distributions, which can hinder fine-detail recovery and lead to suboptimal reconstruction quality.
@@ II-A Single-Modal Point Cloud Completion
The majority of existing point cloud completion methods take partial point clouds as the sole input modality [62, 60, 66]. These works can be broadly categorized into 3D CNN-based methods [6, 56, 67], point-based methods [62, 45, 2, 26, 63, 48], and Transformer-based methods [60, 23, 53, 66].
@@
Early approaches [6, 56, 67] voxelize point clouds into regular grids and apply 3D CNNs in a manner analogous to 2D image processing. While effective at low resolutions, voxel-based representations incur prohibitive computational and memory costs as resolution increases, which severely limits their scalability. Motivated by the success of PointNet [32] and its variants [33, 47, 8], PCN [62] and subsequent point-based methods [45, 2, 26, 63, 48] operate directly on unordered point sets via shared MLPs, typically coupled with folding-based decoders [58]. For example, CRN [45] introduces a two-stage pipeline that refines coarse predictions through a lifting module, while ODGNet [2] employs an orthogonal dictionary to guide training. However, both approaches require learning dataset-specific shape priors, which may restrict generalization to unseen objects. MSN [26] and NSFA [63] formulate point cloud generation as a deformation process, whereas TopNet [38] and PMPNet [48] model the generation process using tree structures and earth mover formulations, respectively. Despite their effectiveness, many of these methods depend on tightly constrained generators with limited representational capacity, which can lead to oversmoothed surfaces and suboptimal reconstruction quality.
@@
Inspired by advances in natural language processing and 2D vision, recent studies have introduced Transformer architectures [41] to capture global dependencies within point clouds. As a pioneering work, PoinTr [60] proposes a geometric-aware Transformer that incorporates KNN operators to exploit local geometric priors. Subsequent works largely follow two main directions: introducing additional inductive biases [59, 5, 12, 50] and designing specialized architectural components [3, 21, 35]. In the first line, PPCC [59] and PaCo [5] investigate completion based on object parts and planar primitives, respectively. TCorresNet [12] and DCPCN [50] resort to external data structures such as spherical templates and interfaces, which often require extensive pre-processing and hyperparameter tuning. In the second line, CRA-PCN [35] and PointAttn [43] study local-global fusion strategies, while SeedFormer [66] and ProxyFormer [23] focus on modifications of Transformer architectures tailored to missing-part sensitivity and upsampling. Despite these efforts, the increasing reliance on complex inductive biases and intricate module designs may limit scalability and generalization under large-scale data regimes [51, 19]. In contrast, we shift our focus to more scalable design based on pure self-attention and cross-attention, combined with large-scale training [27, 42] and further exploit complementary information from multiple modalities to improve completion quality.
@@ II-B Cross-Modal Point Cloud Completion
Cross-modal point cloud completion methods include multi-view approaches [54, 69, 16, 55, 14] and image-assisted approaches [64, 1, 70, 57, 20]. Multi-view methods, such as SVDFormer [69] and AEDNet [16], follow a common paradigm of projecting partial point clouds into depth images from multiple viewpoints, which are then used as auxiliary inputs. SparseNet [55] and MMDR [14] further introduce rendering-based losses by comparing projections of the completed point clouds with Ground-Truth (GT) projections. However, in these approaches, the image modality is derived directly from the input point cloud itself, and thus does not provide additional external information beyond the original geometry.
Another line of research incorporates RGB images to assist point cloud completion [64, 1, 70, 57, 20]. ViPC [64] addresses cross-modal learning by training a generative network to regress complete point clouds from single-view images. Nevertheless, bridging RGB apperance to precise 3D geometry is non-trivial, which may limit ViPC in recovering accurate and complete geometry, especially for occluded regions. Subsequent methods, such as XMFNet [1] and CSDN [70], improve cross-modal fusion in the latent space using cross-attention mechanisms and affine normalization, respectively. EGIINet [57] introduces a feature transfer loss to explicitly guide the modality alignment. More recently, BfR [20] explores a retrieval-based paradigm that leverages similar reference samples from a database. However, its performance heavily depends on the quality and scale of the database and incurs additional computational overhead during inference. In contrast to existing approaches that primarily focus on sophisticated modality fusion strategies, we adopt a unified framework that integrates point clouds, RGB images, and textual descriptions for point cloud completion. Furthermore, the proposed modality dropout mechanism effectively reduces overfitting while improving robustness and flexibility under missing-modality scenarios.
@@ II-C Point Cloud Completion Datasets
Dataset plays a crucial role for learning-based point completion methods [27]. Early efforts leverage 3D models from the ModelNet [52] dataset to evaluate shape completion. PCN [62] subsequently introduces the first widely adopted benchmark based on high-quality ShapeNet [4] models. Later datasets expand upon PCN by incorporating more categories and viewpoints. In particular, ShapeNet-34 and ShapeNet-55 [60] utilize all 55 categories from ShapeNetCore [4] and stratify test samples according to different difficulty levels. However, the crop-based strategy used to generate partial point clouds deviates from realistic sensing processes of depth cameras and laser scanners. The MVP dataset [31] addresses this limitation by rendering partial observations using diverse and uniformly distributed camera poses, although the number of paired samples remains relatively limited. KITTI [17] is frequently employed to evaluate real-world generalization, but it contains only vehicle point clouds and lacks ground-truth complete shapes. More recently, ViPC [64] constructs a large-scale dataset with paired RGB images to facilitate cross-modal learning.
@@ III-D Progressive Decoding
The generator maps fused features back to 3D coordinates, which directly impacts the quality of completed point clouds [68]. Previous methods commonly adopt folding-based or MLP-based generators [62, 60, 61], which either deform a 2D grid into 3D surfaces or regress coordinates in a single step. While effective in simple settings, these generators often have limited generative capacity to represent complex and multimodal geometric distributions, and the one-shot regression paradigm can be sensitive to missing observations and prone to producing uneven point densities. Inspired by point cloud upsampling literature [11, 37], we design a progressive generator that performs coarse-to-fine reconstruction through iterative refinement.
@@ IV-B MGPC-1M Benchmark
As summarized in Table I, the training split of MGPC-1M covers over 1k categories spanning a broad range of daily objects, offering substantially greater category diversity than existing point cloud completion datasets. MGPC-1M contains over one million training pairs, exceeding the combined scale of prior benchmarks in this domain. Unlike most existing datasets that provide only point clouds, MGPC-1M additionally includes RGB images and text descriptions, supporting multimodal evaluation and learning. Beyond scale, MGPC-1M adopts more realistic data generation protocols, including depth back-projection with sensor-like noise and input-centric normalization based on the partial observation, which better matches real-world deployment where ground-truth geometry is unavailable. In contrast to datasets that rely solely on synthetic data, our test split is constructed from the real-world GSO dataset and is entirely disjoint from the training data, enabling rigorous evaluation of cross-domain generalization [19]. The resulting test set contains 1,031 novel instances with high-quality real-world models.
@@ VI Conclusion
In this paper, we present a multimodal framework for generalizable point cloud completion that jointly leverages geometric and semantic cues from point clouds, RGB images, and text within a unified Transformer-based architecture. A modality dropout strategy is introduced to improve robustness under varying modality availability, and a progressive decoder is designed to enable coarse-to-fine reconstruction with improved structural consistency and visual fidelity. To facilitate scalable learning and evaluation, we further develop a VLM-assisted data generation pipeline and construct a large-scale benchmark covering over 1,000 categories and one million training pairs. Extensive experiments on both the benchmark and in-the-wild real-world data demonstrate that the proposed approach consistently improves reconstruction quality and exhibits strong zero-shot generalization, highlighting its effectiveness in enhancing the accuracy and robustness of 3D measurement and reconstruction under real-world sensing conditions.